# Generative Hierarchical Temporal Transformer for Hand Action Recognition and Motion Prediction

### v2

@@ 3.2 Joint Modeling of Recognition and Prediction with Semantic-Temporal Hierarchy
A-Block
models the relationship between the mid-level feature and action: it exploits the long-term time span to aggregate the sequence of mid-level features $\mathbf{m}^{P}$ from the whole observation, and predicts a sequence of mid-level features $\mathbf{m}^{A}$ for future timestamps, which are further expanded by $\mathbf{D}^{P}$ into concrete motion. In addition to variational auto-encoding, $\mathbf{A}$ has its latent bottleneck feature also aligned with text embeddings of the action taxonomy, to enable action recognition of the observation and action controlled prediction.
The encoder $\mathbf{E}^{A}$ takes the sequence of mid-level $\mathbf{m}^{P}_{1:n}$ ($n=\lceil T/t\rceil$) and clip-wise object feature ${\mathbf{\omega}}_{1:n}$ as input, to derive action from hand motion and object feature across the observation. In detail, we denote its input as $(\tilde{\mu}^{A},\tilde{\Sigma}^{A},\mathbf{m}^{P}_{1}+\phi_{1},\mathbf{\omega}_{1}+\phi_{1},...,\mathbf{m}^{P}_{n}+\phi_{n},\mathbf{\omega}_{n}+\phi_{n})$, where $\tilde{\mu}^{A},\tilde{\Sigma}^{A}\in\mathbb{R}^{d}$ are trainable tokens. For each clip $i\in[1,n]$, $\mathbf{m}^{P}_{i}$ is the $\mu^{P}$ from $\mathbf{E}^{P}$; $\mathbf{\omega}_{i}$ is aggregated by an extra individual $\mathbf{E}^{PO}$ from the per-frame object features (Sec. 3.4), and is comparable to the CLIP [6, 38, 41] feature of object name; $\phi_{i}$ is the sinusoidal phase encoding denoting the number of clips since the beginning of the performed action. Given the $\mathbf{E}^{A}$ output $(\mu^{A},\Sigma^{A},\eta_{m,1},\eta_{o,1},...,\eta_{m,n},\eta_{o,n})$, the bottleneck latent feature $\mathbf{\alpha}$ is then re-parameterized by $N(\mu^{A},\Sigma^{A})$.
We inject $\mathbf{\alpha}$ into decoder to enable action controlled generation. Specifically, the parallel decoder $\mathbf{D}^{A}$ takes $\mathbf{\alpha}$ for cross-attention, with $\eta_{m,1:n}$ optionally included as skip-connections for enhanced continuity. $\mathbf{D}^{A}$ takes the phase embeddings $\phi_{n+1:n+\bar{n}}$ as input, and outputs $\mathbf{m}^{A}_{n+1:n+\bar{n}}$ depicting the mid-level features of the future $\bar{n}$ consecutive clips. The predicted $\mathbf{m}^{A}$ can then be further expanded into concrete poses through $\mathbf{D}^{P}$, completing the cycle of long-term observation for long-term prediction, with consistency in both global action and local poses (Sec. 3.3, P.b).
To train $\mathbf{A}$, besides the KL-loss $L_{KL}^{A}$, we constrain $\mathbf{\alpha}$ to match the embedding of action taxonomy $\mathcal{A}=\{a=\mathtt{FC_{1}}(\overline{\mathbf{\alpha}})\in\mathbb{R}^{d}\}$, where $\overline{\mathbf{\alpha}}$ are CLIP text embeddings [6, 38, 41] for action labels in the taxonomy of size $N_{A}$. The action recognition loss is:
$L_{action}=\sum_{i=1}^{N_{A}}w_{i}\left(||\mathbf{\alpha}-a_{i}||_{1}-\log\text{Pr}(a_{i}|\mathbf{\alpha})\right)\vspace{-2mm}$ (4)
@@
$\text{Pr}(a_{i}|\mathbf{\alpha})=\frac{\text{exp}(\hat{\mathbf{\alpha}}\cdot\hat{a}_{i}/\tau)}{\sum_{j=1}^{N_{A}}\text{exp}(\hat{\mathbf{\alpha}}\cdot\hat{a}_{j}/\tau)}\vspace{-1mm}$ (5)
measures the probabilistic similarity of predicted and GT labels among candidates from taxonomy. $\hat{z}=z/||z||$ denotes the normalized unit vector, and $\tau=0.07$ is the temperature of contrastive similarity. When testing, we perform action recognition by searching closest labels to $\mu^{A}$.
For future motion supervision, instead of expanding down to concrete hand poses, we directly compare mid-level features for efficiency. In particular, $\mathbf{m}^{A}_{n+1:n+\bar{n}}$ are compared with pre-computed $\overline{\mathbf{m}}^{P}_{n+1:n+\bar{n}}$, where $\overline{\mathbf{m}}^{P}_{j}=\mu^{P}_{j}$ encodes the poses $\overline{\mathbf{H}}_{T+(j-1)t+1:T+jt}$ of the future $j$-th clip via $\mathbf{E}^{P}$. Therefore, the motion prediction loss is
$L_{mid}=\sum_{j=n+1}^{n+\bar{n}}||\mathbf{m}^{A}_{j}-\overline{\mathbf{m}}^{P}_{j}||_{1}\vspace{-2mm}$ (6)
@@ 3.3 Network Flow for Tasks
The framework addresses diverse tasks of recognition (including pose refinement and action recognition) and prediction by going through different paths within the network.
Recognition
@@
    (1)normal-$\to$(4). It generates locally consistent motions using only $\mathbf{P}$. It takes the last $t$ observed frames $\tilde{\mathbf{H}}_{T-t+1:T}$ as input, and predicts motion of the following clip $\mathbf{H}_{T+1:T+t}$ by sampling from $N(\mu^{P},\Sigma^{P})$ for $\mathbf{D}^{P}$ decoding. The output motion can then be autoregressively fed back to $\mathbf{P}$ as input for longer prediction.
@@
    (1)normal-$\to$(2)normal-$\to$(3)normal-$\to$(4). For more realistic long-term prediction with action guidance, we move up the hierarchy to leverage $\mathbf{A}$ and predict $\mathbf{m}^{A}_{n+1:n+\bar{n}}$, with diversity coming from sampling from $N(\mu^{A},\Sigma^{A})$. $\mathbf{m}^{A}_{n+1:n+\bar{n}}$ are further decoded by $\mathbf{D}^{P}$ into concrete poses $\mathbf{H}_{T+t+1:T+(\bar{n}+1)t}$.
@@ 3.4 Implementation Details
We first train $\mathbf{P}$ on all available pose sequences, regardless of the availability and transition of action labels, thanks to the decoupling of action and $\mathbf{P}$. We augment the input motion $\tilde{\mathbf{H}}$ with random noise, making $\mathbf{P}$ capable to cope with noisy per-frame estimation in the recognition stage. Then, we fix the pre-trained $\mathbf{P}$ and train $\mathbf{A}$, whose training data assumes the same action shared between the observation and prediction. We derive the mid-level $\overline{\mathbf{m}}^{P}$ from the pose annotations with $\mathbf{E}^{P}$, randomly divide the training sequence into observed and predicted parts, and correspondingly assign $\overline{\mathbf{m}}^{P}$ as the input of $\mathbf{E}^{A}$ or supervision signal of $\mathbf{D}^{A}$. For the input of $\mathbf{E}^{A}$, we augment the mid-level features with random Gaussian noise, and refer to the noun of GT action for the clip-wise object feature $\mathbf{\omega}$.
We use AdamW [27] optimizer with a learning rate of $10^{-4}$ and weight decay of 0.01 for both $\mathbf{P}$ and $\mathbf{A}$, using a batch size of 256. We respectively train $\mathbf{P}$ and $\mathbf{A}$ with 80 and 200 epochs, with loss weights as $\lambda_{1},\lambda_{2}=1,\lambda_{3}=10^{-5}$, and $\lambda_{4}=1,\lambda_{5}=0.1,\lambda_{6}=10^{-5}$. Other design and training details are illustrated in the supplementary.
Image-based estimator
$\mathbf{F}$ takes an image as input, and outputs for the image its hand pose, along with the object feature $\tilde{\mathbf{o}}$ that is comparable with CLIP [6, 38, 41] feature of object in manipulation. For experiments, we implement $\mathbf{F}$ as a ResNet-18 [17] backbone followed by heads regressing the hand pose and object feature. We provide more details in the supplementary.
Clip-wise object detector
@@ 4.1 Datasets
To demonstrate the versatility of our framework, we use three large-scale hand action datasets [22, 42, 35], covering highly diverse motions and actions, for training and testing. Across all datasets, we consider $N=20$ joints annotated by [22], leaving the carpometacarpal joint of the thumb out as its annotation is unavailable in [42]. We train $\mathbf{P}$ on untrimmed sequences that could contain multiple action annotations, because the short-span hand motion is independent of action, and train $\mathbf{A}$ on trimmed sequences with clean and complete action annotations. For evaluation of both pose and action, we use trimmed sequences with clean action labels.
H2O [22]
records four subjects performing 36 indoor daily activities, in four fixed and one egocentric camera viewpoints. We evaluate on both the validation and test splits, the latter having subjects unseen in training.
Assembly101 [42] and AssemblyHands [35]
Assembly101 [42] contains procedures of people assembling and disassembling toy vehicles, with pose labels computed automatically by UmeTrack [16]. AssemblyHands [35] further improves the pose annotation quality for a subset of Assembly101 sequences.
We follow the splits of Assembly101 for training and testing, using actions of its fine-grained taxonomy with 1380 different labels. We conduct evaluations on sequences that have more reliable pose annotations from AssemblyHands, and consider six fixed camera views that have no severe hand occlusions (details in supplementary). We focus evaluation on the validation split, as it contains accessible object labels for action recognition and motion prediction (Secs. 4.3 and 4.4), which the test split lacks.
@@ 4.2 Setup and Metrics
Recognition
We take a whole video sequence depicting the process of an action as input. For evaluation of pose estimation, we use metrics of MPJPE-RA (Mean Per Joint Position Error-Root Aligned) and MPJPE-PA (Procrustes analysis of MPJPE). To deal with the ambiguity of different hand scales, we align estimation and GT by equalizing the average length of palm bones. For action recognition, we report the top-1 classification accuracy (Action Acc.).
Prediction
We divide each action video into segments of 16 frames. Given each segment, we use its pose and object annotation as the input observation, and predict the rest sequence until reaching the end of action or the maximum duration of 96 frames. We generate 20 random samples from $N(\mu,5\Sigma)$ as a trade-off between accuracy and diversity (Secs. 4.3 and 4.4). For evaluation of generative results, we mainly use the widely adopted FID (Frechet Inception Distance) [18] to assess quality, which computes the distributional distance of features between generated and GT motion sequences. The features are obtained from the last layer of a pre-trained transformer-based action recognition network, and the GT sequences are from the evaluation split unseen in training. Smaller FID means more faithful generation. In addition, to explicitly measure generation diversity, we report APD (Average Pairwise Diversity) [55] that computes the average distance between all pairs of 20 generated samples. Larger APD means more diverse generation. More details about the setup and metrics are given in the supplementary.
@@ 4.3 Joint Modeling of Recognition and Prediction
We first demonstrate the enhanced capability because of our joint modeling of recognition and prediction, by respectively comparing with state-of-the-art solutions for either recognition (i.e. HTT [52]) or prediction (i.e. Pose-GPT [28]). More implementation details for the baselines are given in the supplementary.
Resnet-18($\mathbf{F}$) HTT [52] Ours
@@
Table 1: Pose estimation and action recognition results on H2O [22] and AssemblyHands [35]. For hand pose estimation we report MPJPE-RA/-PA in $mm$ for (left,right) hand respectively, where *** denotes training views leveraged for HTT [52]. Please refer to the supplementary for complete results on all camera views, and comparison on the H2O-Val.
Figure 3: Qualitative comparison of pose estimation for HTT [52] and our G-HTT, on camera view v1 of Assembly datasets [35, 42]. More cases on H2O and AssemblyHands datasets are provided in the supplementary.
Recognition
@@
As shown in Tab. 1 and Fig. 3, G-HTT demonstrates robust accuracy on various camera views, for refining the local pose and action recognition, even though G-HTT is never trained on $\mathbf{F}$. In comparison, although HTT fits better on views that are trained on or close to trained ones (e.g., cam2 of H2O), its performance significantly degrades on the other views, even worse than its input obtained from $\mathbf{F}$. The results show that our simultaneous modeling of both recognition and prediction enhances generalization by learning regular motion priors across tasks. In contrast, a recognition-only network is more likely to overfit to particular data distributions.
Prediction
We take PoseGPT [28], a state-of-the-art model for motion prediction with prescribed action, as a baseline for performance evaluation. While the original PoseGPT trains on body motion, we retrain PoseGPT with its official code by combining the three hand pose-action datasets of Assembly101 [42], AssemblyHands [35] and H2O [22] as we do. We evaluate on action sequences longer than 1 sec to better show the differences of prediction.

### v3

@@ 3.1.1 P-Block
3.1.2 A-Block
models the relationship between the mid-level feature and action: it exploits the long-term time span to aggregate the sequence of mid-level features $\mathbf{m}^{P}$ from the whole observation, and predicts a sequence of mid-level features $\mathbf{m}^{A}$ for future timestamps, which are further expanded by $\mathbf{D}^{P}$ into concrete motion. In addition to variational auto-encoding, $\mathbf{A}$ has its latent bottleneck feature also aligned with text embeddings of the action taxonomy, to enable action recognition of the observation and action-controlled prediction.
The encoder $\mathbf{E}^{A}$ derives action from hand motion and object features across the observation. Its input sequence concatenates the trainable tokens $\tilde{\mu}^{A},\tilde{\Sigma}^{A}\in\mathbb{R}^{d}$ with the clip-wise mid-level $\mathbf{m}^{P}_{1:n}$ and object feature ${\mathbf{\omega}}_{1:n}$ ($n=\lceil T/t\rceil$). Specifically, $\mathbf{m}^{P}_{i}$ is the $\mu^{P}$ from $\mathbf{E}^{P}$; $\mathbf{\omega}_{i}$ is comparable to the CLIP [6, 41, 44] feature of object name, which is aggregated by an extra individual $\mathbf{E}^{PO}$ from the per-frame object features (Sec. 3.3). We further add a sinusoidal phase encoding $\phi_{i}$ to $\mathbf{m}^{P}_{i}$ and $\mathbf{\omega}_{i}$, which denotes the number of clips since the beginning of the performed action. Given $\mu^{A},\Sigma^{A}$ output from $\mathbf{E}^{A}$, we follow $N(\mu^{A},\Sigma^{A})$ to re-parameterize and obtain the bottleneck latent feature $\mathbf{\alpha}$.
We then inject $\mathbf{\alpha}$ into the decoder $\mathbf{D}^{A}$ to enable action-controlled generation. For the cross attention of $\mathbf{D}^{A}$, we utilize $\mathbf{\alpha}$ and optionally include the clip-wise feature obtained from $\mathbf{E}^{A}$ for enhanced continuity. The parallel decoder $\mathbf{D}^{A}$ takes the phase embeddings $\phi_{n+1:n+\bar{n}}$ as input, and outputs $\mathbf{m}^{A}_{n+1:n+\bar{n}}$ depicting the mid-level features of the future $\bar{n}$ consecutive clips. One can further expand the predicted $\mathbf{m}^{A}$ into concrete poses through $\mathbf{D}^{P}$, completing the cycle of long-term observation for long-term prediction, with consistency in both global action and local poses (Sec. 3.2, P.b).
To train $\mathbf{A}$, besides the KL-loss $L_{KL}^{A}$, we constrain the bottleneck latent $\mathbf{\alpha}$ by matching it with the embedding of action taxonomy $\mathcal{A}=\{a=\mathtt{FC_{1}}(\overline{\mathbf{\alpha}})\in\mathbb{R}^{d}\}$, where $\overline{\mathbf{\alpha}}$ are CLIP text embeddings [6, 41, 44] for action labels in the taxonomy of size $N_{A}$. Therefore, the action recognition loss is:
$L_{action}=\sum_{i=1}^{N_{A}}w_{i}\left(||\mathbf{\alpha}-a_{i}||_{1}-\log\text{Pr}(a_{i}|\mathbf{\alpha})\right)$ (3)
@@ 3.1.2 A-Block
$\text{Pr}(a_{i}|\mathbf{\alpha})=\frac{\text{exp}(\hat{\mathbf{\alpha}}\cdot\hat{a}_{i}/\tau)}{\sum_{j=1}^{N_{A}}\text{exp}(\hat{\mathbf{\alpha}}\cdot\hat{a}_{j}/\tau)}$ (4)
measures the probabilistic similarity of predicted and GT labels among candidates from taxonomy. $\hat{z}=z/||z||$ denotes the normalized unit vector, and $\tau=0.07$ is the temperature of contrastive similarity. When testing, we perform action recognition by searching the closest labels to $\mu^{A}$.
For future motion supervision, instead of expanding down to concrete hand poses, we directly compare mid-level features for efficiency. Specifically, $\mathbf{m}^{A}_{n+1:n+\bar{n}}$ are compared with pre-computed $\overline{\mathbf{m}}^{P}_{n+1:n+\bar{n}}$, with $\overline{\mathbf{m}}^{P}_{j}=\mu^{P}_{j}$ encoding the GT hand motion of the future $j$-th clip via $\mathbf{E}^{P}$. The motion prediction loss is
$L_{mid}=\sum_{j=n+1}^{n+\bar{n}}||\mathbf{m}^{A}_{j}-\overline{\mathbf{m}}^{P}_{j}||_{1}$ (5)
@@ 3.2 Network Flow for Tasks
The framework addresses tasks of recognition (i.e. pose refinement and action recognition) and prediction by going through different paths within the network.
@@
    (1$\to$(4). It generates locally consistent motions using only $\mathbf{P}$. It takes the last $t$ observed frames $\tilde{\mathbf{H}}_{T-t+1:T}$ as input, and predicts motion of the following clip $\mathbf{H}_{T+1:T+t}$ by sampling from $N(\mu^{P},\Sigma^{P})$ for $\mathbf{D}^{P}$ decoding. The output motion can then be autoregressively fed back to $\mathbf{P}$ as input for longer prediction.
@@
    (1$\to$(2$\to$(3$\to$(4). For more realistic long-term prediction with action guidance, we move up the hierarchy to leverage $\mathbf{A}$ and predict $\mathbf{m}^{A}_{n+1:n+\bar{n}}$, with diversity coming from sampling from $N(\mu^{A},\Sigma^{A})$. $\mathbf{m}^{A}_{n+1:n+\bar{n}}$ are further decoded by $\mathbf{D}^{P}$ into concrete poses $\mathbf{H}_{T+t+1:T+(\bar{n}+1)t}$.
@@ 3.3.1 G-HTT Details
We first train $\mathbf{P}$ on all available pose sequences, regardless of the availability and transition of action labels, thanks to the decoupling of action and $\mathbf{P}$. We augment the input motion $\tilde{\mathbf{H}}$ with random noise, making $\mathbf{P}$ capable of coping with noisy per-frame estimation in the recognition stage. Then, we fix the pre-trained $\mathbf{P}$ and train $\mathbf{A}$, whose training data assumes the same action shared between the observation and prediction. We derive the mid-level $\overline{\mathbf{m}}^{P}$ from the pose annotations with $\mathbf{E}^{P}$, randomly divide the training sequence into observed and predicted parts, and correspondingly assign $\overline{\mathbf{m}}^{P}$ as the input of $\mathbf{E}^{A}$ or supervision signal of $\mathbf{D}^{A}$. For the input of $\mathbf{E}^{A}$, we augment the mid-level features with random Gaussian noise, and refer to the noun of GT action for the clip-wise object feature $\mathbf{\omega}$.
We use AdamW [30] optimizer with a learning rate of $10^{-4}$ and weight decay of 0.01 for both $\mathbf{P}$ and $\mathbf{A}$, where our batch size is 256. We respectively train $\mathbf{P}$ and $\mathbf{A}$ with 80 and 200 epochs, with loss weights as $\lambda_{1},\lambda_{2}=1,\lambda_{3}=10^{-5}$, and $\lambda_{4}=1,\lambda_{5}=0.1,\lambda_{6}=10^{-5}$. Other design and training details are illustrated in the supplementary.
3.3.2 Image-based Estimator
$\mathbf{F}$ takes an image as input, and outputs for the image its hand pose, along with the object feature $\tilde{\mathbf{o}}$ that is comparable with CLIP [6, 41, 44] feature of the object in manipulation. For experiments, we implement $\mathbf{F}$ as a ResNet-18 [18] backbone followed by heads regressing the hand pose and object feature. We provide more details in the supplementary.
3.3.3 Clip-wise Object Detector
@@ 4.1 Datasets
We use three hand action datasets [24, 45, 38] covering highly diverse motions and actions for training and testing. Across all datasets, we consider $N=20$ joints annotated by [24], leaving the carpometacarpal joint of the thumb out as its annotation is unavailable in [45]. As short-span hand motions are independent of action, we train $\mathbf{P}$ on untrimmed sequences that could contain multiple action annotations, and train $\mathbf{A}$ on trimmed sequences with clean and complete action annotations. We use trimmed sequences with clean action labels to evaluate both pose and action.
H2O [24] records four subjects performing 36 indoor daily activities, in four fixed (cam0-3) and one egocentric camera (cam4) viewpoints. We conduct evaluation on validation and test splits, the latter having subjects unseen in training.
Assembly101 [45] and AssemblyHands [38] Assembly101 [45] contains procedures of people assembling and disassembling toy vehicles, with pose labels computed automatically by UmeTrack [17]. AssemblyHands [38] further improves the pose annotation quality for a subset of Assembly101 sequences. We follow the splits of Assembly101 for training and testing, using actions of its fine-grained taxonomy with 1380 different labels. We conduct evaluations on sequences that have more reliable pose annotations from AssemblyHands, and consider six fixed camera views that have no severe hand occlusions (details in supplementary). We focus evaluation on the validation split, as it contains accessible object labels for action recognition and motion prediction (Secs. 4.3 and 4.4), which the test split lacks.
@@ 4.2 Setup and Metrics
Recognition We take a whole video sequence depicting the process of an action as input. For pose estimation, we use metrics of MPJPE-RA (Mean Per Joint Position Error-Root Aligned) and MPJPE-PA (Procrustes Analysis of MPJPE) (unit: mm). To deal with the ambiguity of different hand scales, we align estimation and GT by equalizing the average length of palm bones. For action recognition, we report the top-1 classification accuracy (Action Acc.).
Prediction We divide each action video into segments of 16 frames. Given each segment, we use its pose and object annotation as the input observation, and predict the rest sequence until reaching the end of action or the maximum duration of 96 frames. We generate 20 random samples from $N(\mu,5\Sigma)$ as a trade-off between accuracy and diversity (Secs. 4.3 and 4.4). For the evaluation of generative results, we mainly use the widely adopted FID (Frechet Inception Distance) [20] to assess quality, which computes the distributional distance of features between generated and GT motion sequences. The features are obtained from the last layer of a pre-trained transformer-based action recognition network, and the GT sequences are from the evaluation split unseen in training. A smaller FID means a more faithful generation. In addition, to explicitly measure generation diversity, we report APD (Average Pairwise Diversity) [62] (unit: mm) that computes the average distance between all pairs of 20 generated samples. A larger APD means a more diverse generation. More details about the setup and metrics are given in the supplementary.
@@ 4.3 Joint Modeling of Recognition and Prediction
We first demonstrate the enhanced capability because of our joint modeling of recognition and prediction, by respectively comparing with state-of-the-art solutions for either recognition (i.e. HTT [58]) or prediction (i.e. PoseGPT [31]). More implementation details for the baselines are given in the supplementary.
Table 1: Pose estimation and action recognition results. For hand pose estimation we report MPJPE-RA/-PA in $mm$ for (left,right) hand respectively, where * denotes training views leveraged for HTT [58]. Please refer to the supplementary for complete results on all camera views, and comparison on the H2O-Val.
H2O-Test Resnet-18($\mathbf{F}$) HTT [58] Ours
@@
Figure 3: Qualitative comparison of pose estimation for HTT [58] and ours, on camera view v1 of Assembly datasets [38, 45]. More cases are provided in the supplementary.
4.3.1 Recognition
@@
As shown in Tab. 1 and Fig. 3, G-HTT demonstrates robust accuracy on various camera views, for refining the local pose and action recognition, even though G-HTT is never trained on $\mathbf{F}$. In comparison, although HTT fits better on views that are trained on or close to trained ones (e.g., cam2 of H2O), its performance significantly degrades on the other views, even worse than its input obtained from $\mathbf{F}$. The results show that our simultaneous modeling of both recognition and prediction enhances generalization by learning regular motion priors across tasks. In contrast, a recognition-only network is more likely to overfit particular data distributions.
Figure 4: Qualitative comparison of predicted motions for PoseGPT [31], the ablated settings of w/o mid-level, w/ only $\mathbf{P}$ via path P.a, and the full G-HTT (w/ $\mathbf{P},\mathbf{A}$, via path P.b) on H2O. More qualitative cases are provided in the supplementary.
GT action input H2O-val H2O-Test AssemblyHands-Val
[…]
PoseGPT [31] $\checkmark$ 5.19 32.3,43.1 11.70 24.1,48.6 16.07 25.3,33.0
Ours (P.b, w/ $\mathbf{P},\mathbf{A}$) $\times$ 5.32 22.1,25.7 8.19 20.1,33.9 5.04 28.1,32.8
4.3.2 Prediction
We take PoseGPT [31], a state-of-the-art model for motion prediction with prescribed action, as a baseline for performance evaluation. While the original PoseGPT trains on body motion, we retrain PoseGPT with its official code by combining the three hand pose-action datasets of Assembly101, AssemblyHands and H2O as we do. We evaluate on action sequences longer than 1 sec to better show the differences in prediction.